Carla is now twice as old as Jennifer. Seven years ago, the sum of their ages was 16. How old is each girl now?
Answer by mananth(16946)

let jennifer's age be x
carla's age = 2x
7 years ago
jennifer was x-7 years
Carla was 2x-7 years
x-7 + 2x-7 =16
3x-14=16
3x=16+14
3x=30
x=10 which is jennifer's age now
Carla's age will be 2* 10 = 20 years